# AgileFormer: Spatially Agile Transformer UNet for Medical Image Segmentation

### v1

@@ 2.2.2 Neighborhood Multi-head Self-Attention (NMSA)
In contrast to standard self-attention, which computes the similarity of each element at a given position $p$ on a feature map $\boldsymbol{f}$ with every other element, the construction of neighborhood attention [23] only leverages the information from the $k$-nearest neighbors around location $p$. We reuse the notation $k$ in Equation˜2, as the NMSA operates like a convolution. As a result, neighborhood attention reduces the quadratic computational complexity in standard self-attention to approximately linear w.r.t. the spatial dimension of $\boldsymbol{f}$, as $k$ is typically small (e.g., $k=7\times 7\ (\times 7)$). Furthermore, this reintroduces local operations into self-attention, allowing for translational equivariance, and thereby enhancing the ability to better preserve local information. Following the notation in Equation˜2, the neighborhood multi-head attention at position $p_{l}$ is computed as
@@
where:
$\begin{split}\hat{\boldsymbol{K}}_{h}[p_{l}]&=[\boldsymbol{K}_{h}[p_{l1}],\cdots,\boldsymbol{K}_{h}[p_{lk}]],\quad\boldsymbol{K}_{h}=\boldsymbol{f}\boldsymbol{W}_{h}^{K}\\ \hat{\boldsymbol{V}}_{h}[p_{l}]&=[\boldsymbol{V}_{h}[p_{l1}]^{\top},\cdots,\boldsymbol{V}_{h}[p_{lk}]^{\top}]^{\top},\quad\boldsymbol{V}_{h}=\boldsymbol{f}\boldsymbol{W}_{h}^{V}.\end{split}$ (5)
Here, $p_{lk}]$ denotes the $k$-th neighboring locations for a given location $p_{l}$. It is worth noting that the dimension of the resulting attention is $\mathbb{R}^{L\times K}$ with $K=k\times k(\times k)$, instead of $\mathbb{R}^{L\times L}$ in standard self-attention or window-attention.
2.3 Multi-scale Deformable Positional Encoding
Figure 3: The proposed multi-scale deformable positional encoding for irregularly sampled grids in deformable multi-head self-attention.
The design of positional encoding (PE) is barely explored by previous ViT UNets. Most ViT-UNets either disregard positional encoding [1] or inherit PEs from their ancestor models [25, 2, 16]. Specifically, [25] used an absolute PE (APE), which assigns an absolute value for each token. While the others [2, 16, 20] used a relative PE (RPE) [26] to encode the relative positions between tokens. However, these are designed for 1D signals while neglecting the spatial correlation. Consequently, they are not well adapted for modeling 2D/3D signals with spatial correlations. Recently, conditional PE (CPE) [27] was designed for vision tasks offering the ability of both APE and RPE at the same time. More importantly, standard APE, RPE, and CPE are not directly applicable to irregularly sampled grids, as all of them are formulated in a rigid grid. In accordance with the irregularly sampled DMSA presented in this paper, we proposed a multi-scale deformable positional encoding (MS-DePE) designed to encode irregularly sampled positional information across various scales (see Figure˜3).
The proposed MS-DePE is formulated in a conditional positional encoding fashion [27] as follows:
$\texttt{MS-DePE}(\boldsymbol{f})=\boldsymbol{f}+\texttt{P}_{\theta}(\boldsymbol{f}),$ (6)
where $\boldsymbol{f}$ is the input feature map, and $\texttt{P}_{\theta}$ denotes the trainable positional encoding layer parameterized by $\theta$. $\texttt{P}_{\theta}$ is implemented as multi-scale deformable depth-wise convolutional layers with different kernel sizes (i.e., $3\times 3\ (\times 3)$ and $5\times 5\ (\times 5)$). For this purpose, we first recover the spatial resolution of the feature map $\boldsymbol{f}$ before passing it through $\texttt{P}_{\theta}$. After applying $\texttt{P}_{\theta}$, we flatten the feature map back to its original shape (see Fig. 3 for details).
2.4 Model construction
@@
Similar to [1, 2, 11], we developed two variants of AgileFormer by varying the embedding dimension ($d_{f}$) and the number of heads ($H$), but kept the main structures unchanged (i.e., number of transformer blocks in the encoder/decoder): AgileFormer-T(iny) ($d_{f}=[64,128,256,512]$; $H=[2,4,8,16]$) and AgileFormer-B(ase) ($d_{f}=[128,256,512,1024]$; $H=[4,8,16,32]$). We also incorporated deep supervision (DS) outlined in [7, 20] into the proposed method.
@@ 3.1 Experimental design
Dataset. In line with previous works [1, 2, 20], we validated the effectiveness of the proposed method on three publicly available medical image segmentation datasets: the Synapse dataset [28], the Automated Cardiac Diagnosis Challenge (ACDC) dataset [29], and the brain tumor segmentation dataset from the Decathlon challenge [30]. The Synapse multi-organ dataset includes 30 3D abdominal CT scans with corresponding segmentation masks of 8 organs (i.e., aorta, gallbladder, kidney left, kidney right, liver, pancreas, spleen, stomach). The ACDC dataset consists of 100 3D cardiac MRI scans, each with a segmentation mask that features the right ventricle (RV), myocardium (Myo), and left ventricle (LV). The Decathlon brain tumor dataset comprises 484 3D multi-modal brain tumor MRI scans with segmentation masks delineating enhancing tumor, non-enhancing tumor, and edema. To be consistent with previous works [18, 20], we reported the results of the whole tumor (WT), enhancing tumor (ET) and tumor core (TC).
Experimental setup and evaluation metric. We adhered to the experimental protocols outlined in [1, 2] for the Synapse and ACDC datasets, including training/testing sets, input image size (i.e., $224\times 224$), data augmentations, model selection, and evaluations. For a fair comparison with the previous literature, experiments for Synapse and ACDC datasets were carried out in 2D. For the 3D volumetric brain tumor segmentation, we followed the experimental protocols specified in [18] for a fair comparison with the previous literature. The input image was set to $96\times 96\times 96$. We evaluated the segmentation performance using the dice similarity coefficient (DSC). Statistical significance between the average DSC of the best and second-best model was estimated using the paired t-test. We emphasize that all competing methods applied in the Synapse, ACDC, and Decathlon datasets followed the same experimental protocols.
Table 1: Performance comparison with 2D methods in the multi-organ segmentation using the Synapse dataset. The best result within each column is highlighted by bold, and the second-best is highlighted with an underline. : models implemented by us. The other benchmarks were taken from [1, 2, 11, 19, 16, 31, 32, 33]. (: $p<0.01$; with the paired t-test to the second-best model. : Large models that have a similar parameters to the AgileFormer-B model (TransCASCADE: 123.47M; MERIT: 147.86M; AgileFormer-B: 112.97M), where deep supervision is also applied.)
[…]
CNN
U-Net [3] 76.85 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58
R50 Att-UNet [1] 75.57 55.92 63.91 79.20 72.71 93.56 49.37 87.19 74.95
Att-UNet [4] 77.77 89.55 68.88 77.98 71.11 93.57 58.04 87.30 75.75
Hybrid
TransUNet [1] 77.48 87.23 63.13 81.87 77.02 94.08 55.86 85.08 75.62
MixedUNet [19] 78.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81
CoTr [12] 78.56 87.09 65.37 86.19 80.32 94.22 52.28 87.01 76.00
Hiformer [16] 80.29 85.63 73.29 82.39 64.84 94.22 60.84 91.03 78.07
TransCASCADE [33] 82.68 86.63 68.48 87.66 84.56 94.43 65.33 90.79 83.69
ViT
R50 ViT [1] 71.29 73.73 55.13 75.80 72.20 91.51 45.99 81.99 73.95
SwinUNet [2] 79.13 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
TransDeepLab [34] 80.16 86.04 69.16 84.08 79.88 93.53 61.19 89.00 78.40
SDAUT [13] 80.67 87.03 69.30 81.87 80.20 94.91 64.56 89.54 77.91
CATFormer [35] 82.17 88.98 67.16 85.72 81.69 95.34 66.53 90.74 81.20
MissFormer [31] 81.96 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
MERIT [11] 84.90 87.71 74.40 87.79 84.85 95.26 71.81 92.01 85.38
MERIT 82.89 87.42 72.61 87.94 85.36 94.45 67.63 87.15 80.53
AgileFormer-T w/o DS 82.74 88.08 75.16 82.41 81.36 95.09 67.23 90.94 81.61
AgileFormer-T w/ DS 83.59 88.81 74.43 84.61 82.78 95.48 69.45 90.14 83.05
AgileFormer-B w/o DS 84.14 87.76 74.71 86.69 83.81 95.31 68.28 91.00 85.58
AgileFormer-B w/ DS 85.74 89.11 77.89 88.83 85.00 95.64 71.62 92.20 85.63
AgileFormer-B w/ DS [best] 86.11 89.29 78.51 89.09 85.12 95.96 71.52 93.03 86.35
Note: MERIT uses the testing set to select the best model, which is not fair for comparison. The evaluation of MERIT using the last-epoch checkpoint results in a lower performance. AgileFormer can even achieve higher performance when using the testing set for model selection. We report this result as AgileFormer-B w/s DS [best] only for comparison, but we highly disencourage any reader from citing this result.
Table 2: Performance comparison with 2D methods in cardiac MRI segmentation using the ACDC dataset. : models implemented by us. The other benchmarks were taken from [1, 2, 19, 31, 33, 11].
@@
Table 3: Performance comparison with 3D models in brain tumor segmentation using the Decathlon dataset. : models implemented by us. The input size was set to $96\times 96\times 96$. Please refer to Appendix˜A for detials of implementing the 3D experiments. We only compared the tiny model (AgileFormer-T), as it has a similar scale to the baselines.
@@
Implementation details. All models were trained using a combination of dice and cross-entropy loss [2, 20], employing an AdamW [36] optimizer with cosine learning rate decay and a 20-epoch warmup. We initialized model parameters with ImageNet pre-trained weights. In the 2D experiments, a batch size of 24 was used, the initial learning rate was set to $3\times 10^{-4}$, and the model was trained for 400 epochs. For the 3D experiments, the batch size and learning rate were set to 4 and $5\times 10^{-4}$, and the model was trained for 200 epochs. We would like to point out that we reimplemented all the 3D baseline models using the nnFormer framework (see Appendix˜B for details of 3D experiments). All experiments were implemented using PyTorch and were performed on a Nvidia V100 GPU with 32GB memory.
Figure 4: Visual comparison of (a) multi-organ, (b) cardiac, and (c) brain tumor segmentation across Synapse, ACDC, and Decathlon datasets. AgileFormer shows the best qualitative segmentation quality across all three datasets.
@@ 3.2 Main results
Our proposed AgileFormer outperformed recent state-of-the-art (SOTA) methods in all three datasets (Table 13 and 3). Specifically, the proposed method performed better than the second-best model by 0.84, 0.23, and 0.80 in average DSC (%) for Synapse multi-organ, ACDC cardiac, and brain tumor segmentation tasks, respectively. This improvement was shown to be statistically significant. Specifically, the proposed method achieved an average DSC of 85.74, surpassing the other baseline methods in segmenting all 6 out of 8 organs in the Synapse dataset (Table˜1). For the remaining two organs, the proposed method is the second-best performing method. The proposed method achieved an average DSC of 92.55 in the ACDC dataset, surpassing all the baseline methods in segmenting RV, Myo, and LV (see Table˜3). In the 3D volumetric brain tumor segmentation task, the proposed method achieved an average DSC of 85.7, surpassing recent SOTA ViT-UNet methods (Table˜3). In addition, the proposed method demonstrated the best qualitative segmentation quality across these three public datasets (see  Figure˜4).

### v2

@@ 3.2.2 Neighborhood Multi-head Self-Attention
In contrast to standard self-attention, which computes the similarity of each element at a given position $p$ on a feature map $\boldsymbol{f}$ with every other element, the construction of neighborhood attention [9] only leverages the information from the $k$-nearest neighbors around location $p$. We reuse the notation $k$ in Equation 2, as the NMSA operates like a convolution. As a result, neighborhood attention reduces the quadratic computational complexity in standard self-attention to approximately linear w.r.t. the spatial dimension of $\boldsymbol{f}$, as $k$ is typically small (e.g., $k=7\times 7\ (\times 7)$). Furthermore, this reintroduces local operations into self-attention, allowing for translational equivariance, and thereby enhancing the ability to better preserve local information. Following the notation in Equation 2, the neighborhood multi-head attention at position $p_{l}$ is computed as
$\begin{split}\operatorname{NMSA}_{h}(&\boldsymbol{f}[p_{l}])=\operatorname{softmax}(\boldsymbol{Q}_{h}[p_{l}]\hat{\boldsymbol{K}}_{h}[p_{l}]^{\top}/\sqrt{d_{k}})\hat{\boldsymbol{V}}_{h}[p_{l}]\\ \end{split}$
where:
$\begin{split}\hat{\boldsymbol{K}}_{h}[p_{l}]&=[\boldsymbol{K}_{h}[p_{l1}],\cdots,\boldsymbol{K}_{h}[p_{lk}]],\quad\boldsymbol{K}_{h}=\boldsymbol{f}\boldsymbol{W}_{h}^{K}\\ \hat{\boldsymbol{V}}_{h}[p_{l}]&=[\boldsymbol{V}_{h}[p_{l1}]^{\top},\cdots,\boldsymbol{V}_{h}[p_{lk}]^{\top}]^{\top},\quad\boldsymbol{V}_{h}=\boldsymbol{f}\boldsymbol{W}_{h}^{V}.\end{split}$
Here, $[p_{lk}]$ denotes the $k$-th neighboring locations for a given location $p_{l}$. It is worth noting that the dimension of the resulting attention is $\mathbb{R}^{L\times K}$ with $K=k\times k~{}(\times k)$, instead of $\mathbb{R}^{L\times L}$ in standard self-attention and window-attention.
3.3 Multi-scale Deformable Positional Encoding
Figure 4: The proposed multi-scale deformable positional encoding ($\operatorname{MS-DePE}$) for irregularly sampled grids in deformable multi-head self-attention. For illustrative purposes, we take the 2D model as an example, where 2D deformable depth-wise convolution is used. In the 3D model, the 2D deformable depth-wise convolution should be replaced with its 3D counterpart.
The design of positional encoding (PE) is barely explored by previous ViT UNets. Most ViT-UNets either disregard positional encoding [5] or inherit PEs from their ancestor models [38, 4, 12]. Specifically, [38] used an absolute PE (APE), which assigns an absolute value for each token. While the others [4, 12, 47] used a relative PE (RPE) [34] to encode the relative positions between tokens. However, these are designed for 1D signals while neglecting spatial correlations. Consequently, they are not well adapted for modeling 2D/3D signals with spatial correlations. Recently, conditional PE (CPE) [6] was designed for vision tasks offering the ability of both APE and RPE at the same time. More importantly, standard APE, RPE, and CPE are not directly applicable to irregularly sampled grids, as all of them are formulated in a rigid grid. In accordance with the irregularly sampled DMSA presented in this paper, we propose a multi-scale deformable positional encoding ($\operatorname{MS-DePE}$) designed to encode irregularly sampled positional grids across various scales (see Figure 4).
The proposed $\operatorname{MS-DePE}$ is formulated in a conditional positional encoding fashion [6] as follows:
$\operatorname{MS-DePE}(\boldsymbol{f})=\boldsymbol{f}+\operatorname{P}_{\theta}(\boldsymbol{f}),$
where $\boldsymbol{f}$ is the input feature map, and $\texttt{P}_{\theta}$ denotes the trainable positional encoding layer parameterized by $\theta$. $\texttt{P}_{\theta}$ is implemented as multi-scale deformable depth-wise convolutional layers with different kernel sizes (i.e., 3×3(×3)3$\times$ 3\ ($\times$ 3)3 × 3 ( × 3 ) and 5×5(×5)5$\times$ 5\ ($\times$ 5)5 × 5 ( × 5 )). For this purpose, we first recover the spatial resolution of the feature map $\boldsymbol{f}$ before passing it through $\texttt{P}_{\theta}$. After applying $\texttt{P}_{\theta}$, we reshape the feature map back to its original shape (see Fig. 4 for details).
Table 1: Performance comparison with 2D and 3D segmentation networks in the multi-organ segmentation using the Synapse multi-organ segmentation dataset. The best result for 2D and 3D models within each column is highlighted by bold, and the second-best is highlighted with an underline. : models implemented by us. The other benchmarks were taken from [5, 4, 29, 37, 12, 16, 42, 27, 47].
@@ 3.3 Multi-scale Deformable Positional Encoding
U-Net [30] 76.85 39.70 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58
Att-UNet [26] 77.77 36.02 89.55 68.88 77.98 71.11 93.57 58.04 87.30 75.75
TransUNet [5] 77.48 31.69 87.23 63.13 81.87 77.02 94.08 55.86 85.08 75.62
MixedUNet [37] 78.59 26.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81
CoTr [41] 78.56 24.05 87.09 65.37 86.19 80.32 94.22 52.28 87.01 76.00
Hiformer [12] 80.69 19.14 87.03 68.61 84.23 78.37 94.07 60.77 90.44 82.03
PVT-GCASCADE [28] 83.28 15.83 88.05 74.81 88.02 84.83 95.38 69.73 91.92 83.63
Trans-CASCADE [27] 82.68 17.34 86.63 68.48 87.66 84.56 94.43 65.33 90.79 83.69
2D Models R50 ViT [5] 71.29 32.87 73.73 55.13 75.80 72.20 91.51 45.99 81.99 73.95
SwinUNet [4] 79.13 21.55 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
TransDeepLab [2] 80.16 21.25 86.04 69.16 84.08 79.88 93.53 61.19 89.00 78.40
SDAUT [15] 80.67 25.59 87.03 69.30 81.87 80.20 94.91 64.56 89.54 77.91
CATFormer [44] 82.17 16.20 88.98 67.16 85.72 81.69 95.34 66.53 90.74 81.20
MissFormer [16] 81.96 18.20 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
MERIT [29] 84.90 13.22 87.71 74.40 87.79 84.85 95.26 71.81 92.01 85.38
AgileFormer-T w/o DS 82.74 19.28 88.08 75.16 82.41 81.36 95.09 67.23 90.94 81.61
AgileFormer-T w/ DS 83.59 15.09 88.81 74.43 84.61 82.78 95.48 69.45 90.14 83.05
AgileFormer-B w/o DS 84.14 15.27 87.76 74.71 86.69 83.81 95.31 68.28 91.00 85.58
AgileFormer-B w/ DS 85.74 18.70 89.11 77.89 88.83 85.00 95.64 71.62 92.20 85.63
3D Models nnUNet [17] 86.99 10.78 93.01 71.77 85.57 88.18 97.23 83.01 91.86 85.26
UNETR [11] 79.56 22.97 89.99 60.56 85.66 84.80 94.46 59.25 87.81 73.99
SwinUNETR [10] 83.48 10.55 91.12 66.54 86.99 86.26 95.72 68.80 95.37 77.01
nnFormer [47] 86.57 10.63 92.04 70.17 86.57 86.25 96.84 83.35 90.51 86.83
AgileFormer-T w/ DS 87.43 7.81 91.80 73.02 87.63 87.49 97.06 82.05 95.88 84.48
  • : $p<0.05$; with Wilcoxon signed rank test to the second-best model. : Large models that have similar parameters to the AgileFormer-B model (Figure 1 (Right)), where
    deep supervision is also applied. Note: All 2D AgileFormer results reported in this paper use the checkpoint from the last training epoch. If using the testing set to select the
    best model (e.g., as done in the code of MERIT, Trans-CASCADE, and PVT-GCASCADE), 2D Agileformer-B can achieve a DSC (%) and HD95 of 86.11 and 12.88.
3.4 Model Configurations
@@
2D AgileFormer. Similar to [5, 4, 29], we developed two variants of AgileFormer by varying the embedding dimension ($d_{f}$) and the number of heads ($H$), while keeping the main structures unchanged (i.e., number of transformer blocks in the encoder/decoder): AgileFormer-T(iny) ($d_{f}=[64,128,256,512]$; $H=[2,4,8,16]$) and AgileFormer-B(ase) ($d_{f}=[128,256,512,1024]$; $H=[4,8,16,32]$). We also incorporated deep supervision (DS) outlined in [18, 47] into the proposed method.
[…]
@@ 4.1 Experimental design
Dataset. In line with previous works [5, 4, 47], we validated the proposed method on three publicly available medical image segmentation datasets: the Synapse multi-organ dataset [19], the Automated Cardiac Diagnosis Challenge (ACDC) dataset [3], and the brain tumor segmentation dataset from the Decathlon challenge [1]. The Synapse multi-organ dataset includes 30 3D abdominal CT scans with corresponding segmentation masks of 13 organs. For brevity, we refer to this dataset as Synapse hereafter. As consistent with prior works, we report the performance on 8 abdominal organs (i.e., aorta, gallbladder, kidney left, kidney right, liver, pancreas, spleen, stomach). The ACDC dataset consists of 100 3D cardiac MRI scans, each with a segmentation mask that features the right ventricle (RV), myocardium (Myo), and left ventricle (LV). The Decathlon brain tumor dataset comprises 484 3D multi-modal brain tumor MRI scans with segmentation masks delineating enhancing tumor, non-enhancing tumor, and edema. To be consistent with previous works [11, 47], we reported the results of the whole tumor (WT), enhancing tumor (ET) and tumor core (TC).
Table 2: Performance comparison with 2D and 3D segmentation models in cardiac MRI segmentation using the ACDC dataset. : models implemented by us. The other benchmarks were taken from [5, 4, 37, 16, 27, 29, 47].
@@
3D Models UNETR [11] 88.61 85.29 86.52 94.02
UNETR++ [33] 92.83 91.89 90.61 96.00
PHTrans 91.79 90.13 89.48 95.76
nnUNet [17] 91.61 90.24 89.28 95.36
nnFormer [47] 91.84 90.91 89.44 95.17
AgileFormer-T w/ DS 92.07 90.59 89.80 95.81
  • : $p<0.05$; with the paired t-test to the second-best model. We reproduce the
    results of nnFormer by taking the model weights released by the authors in [47].
Table 3: Performance comparison with 3D segmentation models in brain tumor segmentation using the Decathlon dataset. : models implemented by us.
@@
  • : $p<0.05$; with the paired t-test to the second-best model.
Experimental setup and evaluation metric. We evaluated the performance of both 2D AgileFormer (trained using slices) and 3D AgileFormer (trained using volumes), as previous works involved both 2D and 3D models. For 2D AgileFormer, we adhered to the experimental protocols outlined in [5, 4] for the Synapse and ACDC datasets, including training/testing partitioning, input image size (i.e., $224\times 224$), data augmentations, model selection, and evaluations. For the 3D AgileFormer, we followed the experimental protocols specified in [47] for a fair comparison with the previous literature. The input image sizes for Synapse, ACDC, and Decathlon brain tumor datasets were set to $64\times 128\times 128$, $14\times 160\times 160$, and $96\times 96\times 96$  [^1] We used an input size of $96\times 96\times 96$ instead of $128\times 128\times 128$ in nnFormer [47] for brain tumor segmentation. This is because most baseline methods [5, 11, 10, 38, 41] in nnFormer [47] were implemented in $96\times 96\times 96$, and hence show a huge performance difference., respectively. We evaluated the segmentation performance using the dice similarity coefficient (DSC) and 95% Hausdorff Distance (HD95). Statistical significance between the average DSC of the best and second-best model was estimated using the paired t-test. We emphasize that all competing methods applied on the Synapse, ACDC, and Decathlon datasets followed the same experimental protocols.
Implementation details. All 2D and 3D models were trained using a combination of DSC and cross-entropy loss [4, 47], employing an AdamW [25] optimizer with cosine learning rate decay. We initialized model parameters with ImageNet pre-trained weights. Since different hyperparameter settings were used for 2D and 3D experiments, we kindly direct readers to Appendix B for the detailed model and hyperparameter configurations for each task. All experiments were implemented using PyTorch and were performed on a Nvidia V100 GPU with 32GB memory.
@@ 4.2 Main results
Results of 2D Models. 2D AgileFormer performed better than the second-best 2D models (i.e., MERIT [29] for both tasks) by 0.84 and 0.23, in average DSC (%) on the Synapse and ACDC datasets, respectively. This improvement was shown to be statistically significant. Specifically, the proposed method achieved an average DSC (%) of 85.74, surpassing the other baseline methods in segmenting 6 out of 8 organs on the Synapse dataset (see Table 1; 2D Models). For the remaining two organs, the proposed method was ranked the second best. One the ACDC dataset, AgileFormer achieved an average DSC (%) of 92.55, surpassing all the baseline methods in segmenting RV, Myo, and LV (see Table 2; 2D Models). Qualitative results also supported the superiority of AgileFormer. While AgileFormer accurately segmented structures of interest, other methods either under- or over-segmented certain regions (see  Figure 6 (a) and (b)).
Results of 3D Models. When comparing 3D models, 3D AgileFormer also outperformed other state-of-the-art segmentation methods across all three segmentation tasks. Specifically, AgileFormer achieved an average DSC (%) of 87.43 and an average HD95 (mm) of 7.81, surpassing nnFormer by 0.86 and 2.82 in DSC and HD95, respectively (Table 1; 3D Models). On the ACDC dataset, AgileFormer achieved an average DSC (%) of 92.07, outperforming nnFormer by 0.23 (Table 2; 3D Models). On the brain tumor segmentation task, the proposed method achieved an average DSC (%) of 85.7, surpassing nnFormer by 0.8 (Table 3). In addition, AgileFormer showed high qualitative segmentation quality on 3D segmentation tasks (see Figure 1 and Figure 6 (c)). In summary, the proposed AgileFormer was more effective on multi-organ and brain tumor segmentation tasks, where target objects exhibit more heterogeneous appearances than those in cardiac segmentation.
Additional qualitative results on Synapse produced by 3D models are shown in Figure S2, where AgileFormer showcased accurate segmentation compared to other state-of-the-art methods. Some failure cases on Synapse are shown in Figure S3, where even methods that can extract spatially dynamic features (including AgileFormer) cannot accurately segment highly irregular organs such as Gallbladder and Pancreas. Similarly, Figure S4 shows failure cases on brain tumor segmentation tasks, where tissue classes with highly irregular appearance cannot be accurately segmented by most methods, including AgileFormer.
Figure 5: Comparison of model scalability on the Synapse dataset. The base model is almost four times larger than the tiny model. The proposed AgileFormer demonstrated exceptional scalability when adding parameters compared to other methods.